OpenStudy (anonymous):

what's the vertex of x-1=1/12(y-3)^2?

OpenStudy (campbell_st):

the parabola is almost in this form \[4a(x -h) = (y - k)^2\] this is the general equation for a concave right parabola with a vertex (h, k) multiply both sides of your parabola by 12 and you get \[12(x -1) = (y - 3)^2\] now compare your equation with the standard form above to get the vertex